Find the molar mass of the unknown gas which diffuses

Twice as fast as HI
1
Expert's answer
2020-12-03T14:06:44-0500

Molar mass of HI, M1= 1+127= 128gmol-1

let the molar mass of the unknown gas X be M2

Then from Graham's law of diffusion

Rx/RHI = √(M1/M2)

2/1 = √(128/M2)

2²/1² = 128/M2

4 = 128/M2

M2 = 128/4 = 32gmol-1